OpenStudy (anonymous):

In a 30°- 60°- 90° right triangle, the longer leg is cm. How long are the other two sides of the triangle?

OpenStudy (anonymous):

ratios are \(1:\sqrt{3}:2\)

OpenStudy (anonymous):

problem is there was a copy and paste fail, but if the long side is say \(x\) then the short side is \(\frac{x}{\sqrt{3}}\) and the long side is \(\frac{2}{\sqrt{3}}\)

OpenStudy (anonymous):

6 sq root 3 cm*

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\[\frac{6\sqrt{3}}{\sqrt{3}}=6\] for the short side

OpenStudy (anonymous):

long hypotenuse is \(2\times 6=12\)

OpenStudy (anonymous):

site is really lagging, sorry